OpenStudy (anonymous):

Simplify and select the answer with the appropriate restrictions for the variable 3x^2+9x / x^2-2x-15

OpenStudy (anonymous):

Im not gonna lie. I have no clue how to solve this problem

OpenStudy (anonymous):

\[\frac{ 3x^2+9x }{ x^2-2x-15 } = \frac{ 3x(x+3) }{ (x-5)(x+3) } \] The (x+3)s cancel out and you get \[\frac{ 3x }{ x-5 }\]

OpenStudy (anonymous):

thank you